# The anti-glitching gamma-ray pulsar PSR J1522-5735

### v1

@@ 6 Pulse profile and energy flux
Figure 3: The energy flux from PSR J1522-5735 calculated in 75% overlapping 60-day data segments. The horizontal axis shows the time midpoint of the data segment. The shaded region correspond to one-sigma uncertainties.
In order to search for pulse profile changes induced by the glitches, we performed Monte Carlo simulations similar to those previously described. We used the same data segments but permitted variation in the post-glitch pulse profile template during the runs. By ”phase-folding” photon arrival times at each Monte Carlo step, the average weighted pulse profiles were calculated. We found no significant alteration in the pulse profile shape following either the spin-up or the spin-down glitches (see Figure 4).
Figure 4: Weighted pulse profiles of PSR J1522-5735 before the first (blue line) and after the third (red line) glitch events respectively. The error bars show one-sigma uncertainties. The average template pulse profiles used for timing analyses are shown by black dashed-dotted line. Grey dashed line indicates the background level.
It is also important to search for flux variations related to the glitches. For this purpose the gamma-ray flux from PSR 1522-5735 were monitored by dividing the full data into 60-day-long data segments with 75% overlap. Base on the data from each segment, we constructed the spectral model of the gamma-ray emission of the region around the PSR J1522-5735, as previously described in Section 2. The spectral parameters of the sources were fixed to the global ones due to the reduced exposure. The spectral normalization factors of PSR J1522-5735 and of the sources within $5^{\circ}$ region around its position were set free and then optimized for every time segment with binned likelihood analysis by the gtlike tool. The energy flux calculated for each data segment is presented in Figure 3. As a result, we found no significant variations in the flux observed around the epochs of the glitch events.
@@ 7 Discussion
We carried out a comprehensive gamma-ray timing analysis of the pulsar PSR J1522-5735 using the publicly available Fermi-LAT data spanning from 2008 August 4 to 2024 April 5. The analysis revealed eight glitch events during this period: two spin-up glitches and six anti-glitches. The anti-glitches are well fitted by a model with step-like changes in frequency and frequency time derivative with no post-glitch recovery terms. However, a rapid recovery following anti-glitch within a few days cannot be reliably constrained due to the sparsity of gamma-ray data. On the other hand, the timescales for recoveries following these events are shorter than that of the majority of other glitches seen in rotation-powered pulsars, especially those characterized by a single exponential timescale.
The anti-glitch events were radiatively quiet, exhibiting no significant variations in the shape of the pulse profile or energy flux. This may suggest that an internal mechanism is responsible for these phenomena.
While this study was in the final stage, the paper by Zhou et al. (2024) appeared with similar results. In that work, TOAs fitting was used to ascertain glitch parameters, without employing BIC or alternative measures for model selection. This may potentially clarify the differences in the glitch parameters found in this study and in the referenced paper, as some exceed 2-sigma level. Exploring these differences is beyond the scope of this paper and will be addressed in future studies.
@@ Acknowledgements.
We are indebted to O.E. Kalashev, G.I. Rubtsov and Y.V. Zhezher for numerous inspiring discussions. The work is supported by the Russian Science Foundation grant 17-72-20291. The numerical part of the work is performed at the cluster of the Theoretical Division of INR RAS.

### v2

@@ 6 Pulse profile and energy flux
Figure 3: Weighted pulse profiles of PSR J1522-5735 averaged over Monte Carlo realisations for the timing model obtained using the data before the first (blue line) and after the third (red line) glitch events. The error bars show the standard deviation across the ensemble of realisations. The template pulse profile used for the timing analyses is shown with a black dashed-dotted line. The grey dashed line indicates the background level.
In order to search for pulse profile changes induced by the glitches, we performed Monte Carlo simulations similar to those previously described using the same data segments but allowing the post-glitch pulse profile template to vary during the runs. At each Monte Carlo step, we evaluated the weighted pulse profile by ‘phase-folding’ photon arrival times for a given realisation of model parameters. Finally, the weighted pulse profile was averaged over the realisations. We found no significant alteration in the pulse profile shape following either the spin-up or the spin-down glitches (see Figure 3).
It is also important to search for flux variations related to the glitches. For this purpose, the gamma-ray flux from PSR J1522-5735 were monitored by dividing the full data into 60-day-long data segments with 75% overlap. Based on the data from each segment, we constructed the spectral model of the gamma-ray emission of the region around the PSR J1522-5735 as previously described in Section 2. The spectral parameters of the sources were fixed to the global ones due to the reduced exposure. The spectral normalisation factors of PSR J1522-5735 and of the sources within $5^{\circ}$ of the region around its position were set free and then optimised for every time segment with binned likelihood analysis by the gtlike tool. The energy flux calculated for each data segment is presented in Figure 4. As a result, we found no significant variations in the flux observed around the epochs of the glitch events.
Figure 4: Energy flux from PSR J1522-5735 calculated in 75% overlapping 60-day data segments. The horizontal axis shows the time midpoint of the data segment. The shaded region corresponds to one-sigma uncertainties.
@@ 7 Discussion
We have carried out a comprehensive gamma-ray timing analysis of the pulsar PSR J1522-5735 using the publicly available Fermi-LAT data spanning from 2008 August 4 to 2024 April 5. The analysis revealed eight glitch events during this period: two spin-up glitches and six anti-glitches. The anti-glitches are well fitted by a model with step-like changes in frequency and a frequency time derivative with no post-glitch recovery terms. However, a rapid recovery following anti-glitch within a few days cannot be reliably constrained due to the sparsity of gamma-ray data. On the other hand, recoveries in a few days are shorter than those following the majority of glitches observed in rotation-powered pulsars, especially those characterised by a single exponential timescale.
The detection of six anti-glitches compared to two spin-up glitches in a rotation-powered pulsar clearly has potential importance for the understanding of these phenomena and necessitates a re-evaluation of existing models. Models that explain anti-glitches in magnetars (see, for example, García & Ranea-Sandoval 2015) are probably not applicable to the rotation-powered PSR J1522-3755, which does not possess sufficiently strong magnetic fields. External models such as those associated with strong outflows (Tong 2014) or corona-mass-ejection-like events (Lyutikov 2013) can also explain the spin-down events. However, most existing models predominantly yield to gradual deceleration processes, thereby significantly constraining their capacity to generate sudden spin-down events. Alternatively, Huang & Geng (2014) propose that anti-glitches may arise from the collision of a pulsar with a small celestial body. Such collisions could induce a rapid spin-down of the pulsar’s rotation rate. However, in such a scenario, it is still difficult to explain the frequency of anti-glitch events observed in PSR J1522-5735.
The observation of radiatively quiet anti-glitch events characterised by the absence of significant variations in pulse profile shape and energy flux suggests an internal origin for these phenomena. Kantor & Gusakov (2014) proposed an anti-glitch model by extending the standard internal scenario of Anderson & Itoh (1975). In this case, the glitch is due to momentum transfer between the pulsar’s superfluid and normal components in its inner crust or outer core, depending on where the quantised vortices can pin. They proposed that the superfluid fraction is dependent on the superfluid current and increases as the velocity difference between the superfluid and normal components decreases. In this case, under some condition, the pinned superfluid and the rest of the star can decelerate, while the moment of inertia redistributes through the formation of additional Cooper pairs to satisfy angular momentum conservation, resulting in an anti-glitch. The higher the internal temperature ($\geq 10^{7}$ K) and the rotation frequency lag between the components before vortex unpinning ($\geq 1$ rad/s), the more likely it is that an anti-glitch occurs. PSR J1522-5735, with a characteristic age of $52$ kyr, ranks among the youngest 4% of known pulsars. Its youth implies elevated internal temperatures and a dynamic rotational evolution, both of which favour the development of a substantial rotational lag. These properties align with the proposed prerequisites for the anti-glitch scenario of Kantor & Gusakov (2014).
While this study was in the final stage, the paper by Zhou et al. (2024) appeared with similar results. In that work, TOA fitting was used to ascertain glitch parameters without employing BIC or alternative measures for model selection. This may potentially clarify the differences in the glitch parameters found in this study and in the referenced paper, as some exceed the two-sigma level. Exploring these differences is beyond the scope of this paper and will be addressed in future studies.
@@ Acknowledgements.
We are indebted to M.Yu. Kuznetsov and G.I. Rubtsov for numerous inspiring discussions. The work is supported by the Russian Science Foundation grant 17-72-20291. The numerical part of the work is performed at the cluster of the Theoretical Division of INR RAS.